# Non-Hermitian Fabry–Pérot Resonances

### v1

@@ 3.5. Exact exceptional points from parity-time symmetric radiation conditions
To achieve $\mathcal{PT}$-symmetry at the differential equation level, we choose a spatially reflection symmetric resonator array together with a modified radiation condition known as the perfect transmission condition; see [fleury.sounas.ea2015invisible, PhysRevResearch.3.013223]. The modified boundary conditions are:
@@
In Figure˜5, we show the two closest resonances to $\omega=\pi$ of the dimer system, both under standard outgoing radiation conditions (2.6) and under perfect transmission radiation conditions (3.4), demonstrating the role the radiation conditions play in facilitating this exceptional point.
Figure˜4(c) demonstrates square-root perturbation of the resonant frequencies as the left boundary conditions are varied around the perfect transmission condition. The left boundary condition of the system is parametrised by $\theta_{L}\in[0,2\pi)$ and enforced by requiring,
@@ 4. Non-reciprocal systems
We now turn our attention to Fabry-Pérot resonances in non-reciprocal systems. To that end, we modify ˜2.2 by introducing a non-reciprocal imaginary gauge potential ([ammariMathematicalFoundationsNonHermitian2024, yokomizo]) and obtain the following generalised Strum-Liouville equation:
@@ 4.1. Propagation matrix approach
see [pm1, Section 5]. Moreover, by factoring out the non-reciprocal decay as done in the characterisation (4.9), we can find a similar but symmetric system to $\mathcal{C}^{\gamma}$; see [ammari2025competingedgebulklocalisation].
We also note that, analogously to Lemma˜2.3, a characterisation of higher resonant frequencies for non-reciprocal systems can be derived using the same approach as in [pm2]. Nevertheless, we do not make use of it in the sequel.
@@ 4.2. Symmetrisation by gauge transformation
Following standard Sturm-Liouville theory [zettl2005SturmLiouville, Chapter 3], since the potential $V$ is real-valued, formal symmetry is sufficient to ensure self-adjointness of $\mathcal{S}$. Let $f,g\in\operatorname{dom}(S)$. We compute
@@ 4.3. Non-Hermitian skin effect
We recall the following two results from the Floquet-Bloch theory of periodic operators; see [eastham1973spectral].
@@
This corollary is illustrated in Figure˜6 for the resonant modes of a one-resonator unit cell ($N=1$), repeated $M=20$ times. Indeed, both the subwavelength and non-subwavelength resonant modes $u(x)$ decay exactly proportionally to the characteristic decay $e^{-\Gamma(x)}$, as predicted.
We note that using the approach introduced in [ammari2025competingedgebulklocalisation], we can generalise our proof here for the skin effect in the non-subwavelength regime to aperiodic systems.
@@ 4.4. Damping transition
We demonstrate this transition empirically in Figure˜7 by plotting the symmetrised resonant modes $v(x)$ given by (4.4) for two dimer systems with distinct decay rates $\gamma$. We observe that the resonant modes undergo a transition from oscillatory to non-oscillatory within the resonators, as $\gamma$ is increased while $\omega$ remains roughly constant.

### v2

@@ 3.5. Exact exceptional points from parity-time symmetric radiation conditions
To achieve $\mathcal{PT}$-symmetry at the differential equation level, we choose a spatially reflection symmetric resonator array together with a modified radiation condition known as the perfect transmission condition; see [undefl, undeft]. The modified boundary conditions are:
@@
In Figure˜6, we show the two closest resonances to $\omega=\pi$ of the dimer system, both under standard outgoing radiation conditions (2.6) and under perfect transmission radiation conditions (3.4), demonstrating the role the radiation conditions play in facilitating this exceptional point.
Figure˜5(c) demonstrates square-root perturbation of the resonant frequencies as the left boundary conditions are varied around the perfect transmission condition. The left boundary condition of the system is parametrised by $\theta_{L}\in[0,2\pi)$ and enforced by requiring,
@@ 4. Non-reciprocal systems
We now turn our attention to Fabry-Pérot resonances in non-reciprocal systems. To that end, we modify ˜2.2 by introducing a non-reciprocal imaginary gauge potential ([undef, undefv]) and obtain the following generalised Strum-Liouville equation:
@@ 4.1. Propagation matrix approach
see [undefo, Section 5]. Moreover, by factoring out the non-reciprocal decay as done in the characterisation (4.9), we can find a similar but symmetric system to $\mathcal{C}^{\gamma}$; see [undefd].
We also note that, analogously to Lemma˜2.3, a characterisation of higher resonant frequencies for non-reciprocal systems can be derived using the same approach as in [undefh]. Nevertheless, we do not make use of it in the sequel.
@@ 4.2. Symmetrisation by gauge transformation
Following standard Sturm-Liouville theory [undefw, Chapter 3], since the potential $V$ is real-valued, formal symmetry is sufficient to ensure self-adjointness of $\mathcal{S}$. Let $f,g\in\operatorname{dom}(S)$. We compute
@@ 4.3. Non-Hermitian skin effect
We recall the following two results from the Floquet-Bloch theory of periodic operators; see [undefj].
@@
This corollary is illustrated in Figure˜7 for the resonant modes of a one-resonator unit cell ($N=1$), repeated $M=20$ times. Indeed, both the subwavelength and non-subwavelength resonant modes $u(x)$ decay exactly proportionally to the characteristic decay $e^{-\Gamma(x)}$, as predicted.
We note that using the approach introduced in [undefd], we can generalise our proof here for the skin effect in the non-subwavelength regime to aperiodic systems.
@@ 4.4. Damping transition
We demonstrate this transition empirically in Figure˜8 by plotting the symmetrised resonant modes $v(x)$ given by (4.4) for two dimer systems with distinct decay rates $\gamma$. We observe that the resonant modes undergo a transition from oscillatory to non-oscillatory within the resonators, as $\gamma$ is increased while $\omega$ remains roughly constant.
@@ Acknowledgments
Code availability
The software used to produce the numerical results in this work is openly available at
https://doi.org/10.5281/zenodo.18622435.